# I Wish to Have an Argument: Argumentative Reasoning in Large Language Models

### v1

@@ Abstract
We evaluate the ability of contemporary large language models (LLMs) to perform argumentative reasoning. We frame our experiments in terms of the argument mining (AM) and argument pair extraction (APE) tasks, and evaluate their ability to perform reasoning at increasing levels of abstraction in the input and output representations (e.g., arbitrary label sets, semantic graphs). We find that, although LLMs are able to match or surpass the state-of-the-art in AM and APE, their argumentative reasoning performance is very dependent on the input and output representation. We also find an “exemplar effect”, where too many exemplars increasingly become detrimental for task performance, and about $4-5$ being the optimal amount. Neither result extends to chain-of-thought (CoT) prompting: we find the exemplar effect to be nullified, and our results suggest that CoT allows for better performance under ill-conditioned problems. We hope that the work reported contributes to the improvement of argumentative reasoning in LLMs.
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 (Open AI, 2023) are able to solve a significant amount of natural language processing tasks to a spectacular degree of accuracy. Lately, there has been a push to use these models to automate more complex tasks reliant on deeper comprehension of language, such as data labelling (Cheng, Li, and Bing, 2023; Liu et al., 2023), model evaluation (Zheng et al., 2023), and even paper reviews (Liu and Shah, 2023).
Argumentative reasoning is central to all of these efforts. This type of reasoning encompasses formal and informal logic, and has as a prerequisite a deep understanding of language and the context. Plenty of work has been carried out to evaluate LLM performance in formal logic scenarios; but, in human-centered applications, such as chatbots, formal logic issues are overlooked in favor of defeasible reasoning (Walton, 2008). Hence, to accurately decide whether LLMs are reliable on more complex tasks, such as data annotation and human interaction, we must also evaluate their capabilities in the context of argumentative reasoning.
In this paper we perform one such evaluation for GPT-3 (Brown et al., 2020) and GPT-4. We explore these model’s argumentative reasoning capabilities by altering the input and output representations and subsequently observing their performance in two argumentation-related tasks. We frame our experiments from an argument mining (AM) and argument pair extraction (APE; Cheng et al., 2020) perspective. AM requires the model to be able to recognize spans of a text as belonging to an argument; and APE tasks the model to match arguments from two texts as argument-rebuttal pairs.
For the input and output representation evaluation, we test progressively more abstract reasoning settings: from the concrete setting, where the model is tasked to return matches from a passage (e.g., in AM it would return the spans that correspond to an argument); to the purely symbolic setting, where we require from the model to return the answer with arbitrary symbols (e.g., extract the spans, label them based on predefined rules, and only return the labels), or the input is in an abstract format such as abstract meaning representation (AMR; (Banarescu et al., 2013)) graphs. AMR graphs are known to aid in argument comparison and interpretation (Opitz et al., 2021).
We also factor in the number of exemplars to test if the model’s reasoning capabilities depend on their context length, and chain-of-thought (CoT; (Wei et al., 2022b)) prompting, a prompting paradigm that conditions the model to work step-by-step by following a templatized process (“Let’s think about this step-by-step…”). It is known to provide good results in multiple reasoning tasks (Wei et al., 2022b; Kojima et al., 2022; Suzgun et al., 2023). The diagram below displays our evaluation.
[…]
@@ 1.1 Findings
In terms of raw scoring, we find that GPT-4 is able to reach SOTA performance in APE, and near-SOTA in AM. However, our analysis also shows the following:
@@
    Both LLMs fall short when the input and output representations are not carefully conditioned. Concrete reasoning does well, but low-abstraction (i.e., with “hints” as the input representation) symbolic approaches outperform it.
@@
    The performance with respect to the number of exemplars behaves consistently regardless of setting, representation, or task. The LLMs have best performance when observing between 4 and 5 exemplars. More exemplars (up to 40 in some cases) cause a linear decrease in performance.
@@
    Performance drops noticeably, but remains above random, when the input representations are too abstract (e.g., AMR graphs). This does not hold for purely symbolic output representations (e.g., matrices), where the performance drops to near-zero levels.
@@
    The results above do not extend to CoT. We find that, in line with the literature, CoT outperforms analogous non-CoT approaches. But we find that the distributions output by CoT are similar, which suggests that the effect of the input and output representations changes is mitigated by this technique. The “exemplar effect” mentioned above does not extend to CoT either.
@@ 2.1 Data
Throughout this paper, we utilize the Review-Rebuttal Submission-v2 (RRv2) dataset from Cheng et al. (2020). It is comprised of $4,764$ pairs of review and rebuttal passages related to scientific article submissions. The test set size is $474$ pairs, and the data was human annotated by $5$ annotators following a specific guideline. Each passage is sentence-separated, and normally includes multiple arguments. RRv2 is a comprehensive corpus and focuses on long-distance relationships between statements; in addition to having two tasks using the same data (AM and APE), which are both actively researched in argumentative reasoning.
For AM, each sentence is labelled with a BIO scheme, and the model must retrieve (label) each sentence from the review and rebuttal entries. [^1] In the RRv2 corpus, the BIO tags correspond to the Beginning, Inner, and Outer parts of an argument. In this task, the distinction between review and rebuttal is not relevant and each entry is treated as a separate point in the corpus.
For APE, the arguments within each review-rebuttal pair must now be aligned: every argument made by a reviewer must be mapped, when applicable, to a refutation from the rebuttal. This is normally represented as a binary matrix with overlaps (Bao et al., 2022; Cheng et al., 2021). See Figure 1 for examples of the tasks and its labelling scheme. Prior to use we clean the text from tag and sentence delimiters.
2.2 Models Evaluated
We evaluate GPT-4 and the text-davinci-003 variant of GPT-3. The latter is similar to the original GPT-3 model, but further pretrained with reinforcement learning with human feedback (Ouyang et al., 2022; Christiano et al., 2017) and instruction pretraining (Wei et al., 2022a; Sanh et al., 2022; Ouyang et al., 2022). It is sometimes referred to as GPT-3.51, and has a context length of $4,097$ tokens. [^2] https://platform.openai.com/docs/models
Figure 1: Examples of the (raw) data entries in RRv2. The BIO tag is in parentheses, and highlighted in orange and blue the “B” and “I” labels corresponding to an argument. In AM, review and rebuttal (response) passages are considered independent, and the task is to label all the arguments from each passage. In APE, the task is to match arguments from the review to their corresponding refutations. In the above the first argument is unmatched, and the second (“BERT produces…”) pairs to the first argument from the response.
Few details have been released around the architecture, model size, or pretraining procedure for GPT-4. Its authors consider it marginally better than GPT-3 at some tasks, and better at more complex tasks. They evaluate it in benchmarks specific to GPT-4 that do not explicitly include AM or APE. However, some of these benchmarks have an argumentation component, such as the essay portion of the SAT. [^3] A test for entrance to universities in the United States. The variant of GPT-4 we utilized has a context length of $32,768$ tokens. Throughout our experiments, for both models we set the temperature to $0.8$, set the maximum return tokens based on the task, and left all other parameters as their default. To account for randomness, we report the average of five calls.
@@ 2.3 Prompting
The models are well-known to be sensitive to the prompt’s wording (Lu et al., 2022), so prior to starting the work carried out some small experimentation to find out which phrasing would be best. We always include exemplars of the desired output format, and, when the prompts used $n$ exemplars, we used the first $n$ points from the development set. In the case of AM, we also prompted GPT-4 with and without CoT, as it has been shown to improve model’s reasoning capabilities in various tasks. We wrote all our prompts following the structure from Open AI (2023), as we observed it produced more reliable outputs. For CoT we followed the standard phrasing (“Let’s read line-by-line and solve this step by step”) and indicated which sentence was being read, as well as the rationale. For example, “We now read {SENTENCE}. It follows the previous argument, and hence it is labelled with an ‘I’.” Sample prompts can be found in Figures LABEL:fig:sampleamconcrete and 2.444The full set of the prompts, scoring code, and model responses will be available at https://github.com/adewynter/argumentation-llms We used the Open AI API for all our experiments, and a consumer-grade laptop for the data processing.
2.4 Scoring
@@
2.5 Baselines
[…]
Extract from the passage all the arguments. Give
every argument in its own paragraph.
@@ 2.5 Baselines
    Exemplars, if any, are included in the prompt and follow the output representation.
    Otherwise (in the case of zero-shot), we only leave the specified output representation. More complex prompts, such as the one for APE, simply ask for the pairing (e.g., return all arguments from the response that match those of the review”).
[…]
@@
Label the beginning of every argument with a B”.
Label the rest of the argument with an I”.
Label every line that is not part of an argument
with an O”.
@@
Figure 2: Sample symbolic AM prompt. The model is requested to parse each line and return BIO tags. Exemplars, if any, are included in the prompt, and follow the specified output representation. For zero-shot, we only specify the output representation. In other output representations the model is requested to mark the B-label (e.g, “(15) 16 17”). In APE we request the index-based response with matching arguments (e.g., “15, 16, 17: 9, 10, 18, 19”) and convert it into the binary matrix representation for scoring. We also experimented with requesting the binary matrix outright.
@@ 2.6 Settings
We have named our settings (i.e., representations) as concrete and symbolic, to broadly distinguish the approach taken towards representing the task. The former returns full sentences, while the later encompasses a variety of input and output symbols. This is solely for practical purposes: in reality, symbolic approaches cover a variety of abstractions which may be considered “easier” than the concrete approach, and, strictly speaking, the concrete approach is a symbolic approach (Brachman and Levesque, 2004). Below we describe these settings, and we discuss our distinctions further in 3.3.
2.6.1 Concrete
We instruct the model to return a portion of the text based on the instructions given: in AM, the lines corresponding to an argument, and in APE, the argument pairs themselves. To distinguish the “B” and “I” labels, we enforce a specific return format to work with our parsing code via a special token (|START|). For scoring concrete answers, we always expect an exact match. See Figure LABEL:fig:sampleamconcrete for a basic AM concrete prompt.
2.6.2 Symbolic
The model is instructed to return symbols (labels) based on its judgment of the text. This requires more reasoning steps than the concrete setting, given that the model is solving AM and labelling the span with a pre-specified, arbitrary label set from the instructions. See Figure 2 for an example. In the AM symbolic setting we evaluated two types of labels: the BIO tags and the line indices. For APE we evaluated the line indices and the full matrix representation. For AMR, we used amrlib [^5] https://github.com/bjascob/amrlib and a pre-trained BART-based model (Lewis et al., 2020) included in the library. The AMR graphs were created at the sentence-level.
@@ 3.1 AM/APE: Symbolic and Concrete Reasoning
In our first experiment we evaluated the performance of the models in AM and APE in both the concrete setting and various iterations of symbolic settings. The results for the best-performing prompts and settings are in Table 1, and a description of every symbolic setting in Table 2.
[…]
In AM, the best-performing symbolic setting had line indices included in the input representation, and requested the indices of each argument as the output representation. To handle the conversion to a BIO label set, we instructed the model to return the “B” labels as indices enclosed in parentheses. Not including the indices in the input did not lead to an equivalently good performance.
In APE, requesting a full binary matrix output led to extremely poor performance ($9.88\%F_{01}$, well below the random baseline, and closer to guessing all zeros every time). However, both symbolic and concrete approaches performed well with respect to the existing non-LLM-based baselines. We tested other approaches, such as first extracting the arguments and then matching them, but it did not yield sufficiently good results. Due to token-length and budget limitations, we were unable to test CoT and AMR in APE.
CoT approaches had generally better performance than their non-CoT counterparts, even though they use fewer exemplars due to their higher token usage. The only exception to this observation was the best-performing approach (AM: indices inline and indices as output) where the difference was of $3\%$ points–still above the concrete setting.
[…]
GPT-3 (c) $39.86\pm 0.51$ $18.58\pm 0.70$
GPT-4 (c) $64.51\pm 0.53$ $53.84\pm 0.73$
GPT-3 (s) $62.00\pm 0.32$ $20.15\pm 0.91$
GPT-4 (s) $70.63\pm 0.21$ $49.85\pm 0.96$
@@
Table 1: Results for the AM and APE tasks in the symbolic (s) and concrete (c) settings. We report the best-performing prompt for an average of five calls: in the case of symbolic reasoning, the best performing approach in both tasks had indices inline and indices as the output. We also report MLMC and MRC-APE, the two best-performing, non-LLM-based approaches for RRv2. Overall, GPT-4 had good performance, almost matching the existing baselines in AM and outperforming them in APE. GPT-3 had much lower performance.

### v2

# “I’d Like to Have an Argument, Please”: Argumentative Reasoning in Large Language Models

@@ Abstract
We evaluate two large language models (LLMs) ability to perform argumentative reasoning. We experiment with argument mining (AM) and argument pair extraction (APE), and evaluate the LLMs’ ability to recognize arguments under progressively more abstract input and output (I/O) representations (e.g., arbitrary label sets, graphs, etc.). Unlike the well-known evaluation of prompt phrasings, abstraction evaluation retains the prompt’s phrasing but tests reasoning capabilities. We find that scoring-wise the LLMs match or surpass the SOTA in AM and APE, and under certain I/O abstractions LLMs perform well, even beating chain-of-thought–we call this symbolic prompting. However, statistical analysis on the LLMs outputs when subject to small, yet still human-readable, alterations in the I/O representations (e.g., asking for BIO tags as opposed to line numbers) showed that the models are not performing reasoning. This suggests that LLM applications to some tasks, such as data labelling and paper reviewing, must be done with care.
keywords:
[…]
[innertopmargin=-10pt, innerbottommargin=-10pt, innerleftmargin=3pt, innerrightmargin=-2pt, skipabove=0pt, skipbelow=0pt, backgroundcolor=blue!5, linecolor=black, outerlinewidth=0.01pt, font=, ]promptbox
and
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 [1] have shown to have spectacular accuracy on a variety of tasks. Hence, attempts have been made to automate more complex tasks reliant on argumentative reasoning, such as data labelling [2] and scientific paper reviews [3]. Argumentative reasoning encompasses formal and informal logic, and requires a deep understanding of, and reasoning over, the pragmatic context. Hence, to understand the reliability of LLMs in these tasks we must also evaluate their argumentative reasoning capabilities. This goes beyond determining whether the model can generate relevant responses, and asks if it can robustly reason over the context and solve the task.
We evaluate the argumentative reasoning capabilities of two LLMs, GPT-3 [4] and GPT-4 in two tasks, argument mining (AM) and argument pair extraction (APE [5]). [^1] Prompts, code, and outputs are in https://github.com/adewynter/argumentation-llms We do this by an abstraction evaluation: measuring progressively more abstract input and output (I/O) representations. Unlike prompt-phrasing evaluations, to which LLMs are known to be sensitive [6, 7], our evaluation maintains the task description untouched, and only alters the signature of the data. For example, adding line numbers to the input and requesting the output to be from a specific set (e.g., $\{0,1\}$) is a conceptually minor I/O representation change. This retains the task description, but requires some level of reasoning to return a correct and parseable response given the specified signatures. In this paper we then measure the LLMs’ argumentative reasoning capabilities indirectly, by testing their ability to robustly recognize arguments when altering the I/O.
@@ 1.1 Findings
In terms of raw scoring, we find that GPT-4 is able to reach SOTA performance in APE, and near-SOTA in AM. However, our analysis shows that:
@@
    LLM scoring varies dramatically with the abstraction level, which suggests a lack of comprehension of the task. Note that across our experiments the task description remained fixed, and we only altered the I/O representation.
@@
    Symbolic prompting (that is, low-abstraction, hints-enabled inputs) bests other approaches, including chain-of-thought (CoT) [8].
@@
    CoT approaches are robust to abstraction, but its output distributions over abstractions are similar, which we attribute to its templatized nature.
@@
    LLM scores worsen with more exemplars, indicating poor inductive reasoning capabilities.
We conclude that the LLMs are unable to reason reliably in an argumentative setting.
2 Related Work
We discuss evaluations of LLMs as it pertains to argumentative reasoning. For non-LLM-based approaches see [9]; and for a survey on reasoning in LLMs see [10]. LLMs are relatively new, though they typically outperform non-LLM approaches, for example in AM [11] and argument evaluation [12, 13], in terms of raw scoring. Indeed, LLMs have been observed to generalize to (read: score well in) unseen tasks without training, thus raising the question as to whether they can reason about the prompts; or are just regurgitating their training data or returning semantically-close responses. To some, this generalization is an indication of emergent reasoning capabilities [14, 15]; but it has been posed that with better statistics this evidence of emergence disappears [16]. Some tests, such as GPT-4’s own technical report [1], tout remarkable reasoning capabilities. There is also evidence to the contrary, e.g. in code generation [17], scientific questions [18], first-order logic under fictional worlds [19], and arithmetic [20]; more generally, tasks with significant reasoning depths cause LLMs to fail [21, 22]. It has hence been suggested that LLMs do not actually reason, but rely on heuristics (e.g., semantic similarity) [23]. Remark that these studies are limited to the prompts and versions of the models available then. It was suggested that LLMs are not meant for formal reasoning, and it is better to evaluate them in real-world (informal, inductive) scenarios [22]; yet GPT-3 cannot mimic human-like inductive reasoning [24], or understand the prompts [7]. LLMs have also been found to not be competent in legal reasoning, due to their inability to make good arguments [25]. LLMs may also retrieve dialogue acts, in line with their success as chatbots, but do not understand offers in negotiations (i.e., pragmatics) [26].
@@ 3.1 Data
Throughout this paper, we utilize the Review-Rebuttal Submission-v2 (RRv2) dataset [5]. It is a comprehensive corpus focused on long-distance relationships between statements, and includes both AM and APE. It has $4,764$ ($474$ for test) pairs of review and rebuttal passages related to scientific article submissions. Each passage is sentence-separated, and includes multiple arguments. It is human-labelled. For AM, each sentence is labelled with a BIO tag, [^2] In the RRv2 corpus, the BIO tags correspond to the Beginning, Inner, and Outer parts of an argument. and the model must retrieve (label) each sentence from the review and rebuttal entries. In AM the distinction between review and rebuttal is irrelevant: each entry is treated as a separate point in the corpus. For APE, the task is to align the arguments within each review-rebuttal pair: every argument made by a reviewer must be mapped, when applicable, to a response from the rebuttal. This is normally represented as a binary matrix with overlaps [27, 28]. Prior to use we clean the text from tag and sentence delimiters. See Figure 1 for a sample of the corpus.
Figure 1: Examples of the entries in RRv2. The BIO tag is in parenthesis. Highlighted in orange and blue are the “B” and “I” labels corresponding to an argument. In AM, review and rebuttal (response) passages are independent, and the task is to assign BIO tags to every line. In APE, the task is to match arguments from the review to their corresponding responses. In the above the first argument is unmatched, and the second (“BERT produces…”) pairs to the first argument from the response.
3.2 LLMs Evaluated
We evaluate GPT-4 and the text-davinci-003 variant of GPT-3 (“GPT-3.51”). Both models are autoregressive language models, instruction-pretrained [29, 30] and tuned with reinforcement learning with human feedback [30, 31]. For GPT-4, there are no details released around the architecture, model size, or training data. It is considered better than GPT-3 at more complex tasks [1]. The variant of GPT-4 we used (“GPT-4-0613”) has a context length of $32,768$ tokens; GPT-3 has $4,097$ tokens. [^3] https://platform.openai.com/docs/models. Throughout our experiments, for both LLMs we set the temperature to $0.8$, the maximum return tokens based on the task, and left everything else as default. To account for randomness, we report the average of five calls per point to the Azure OpenAI API.
@@ 3.3 Prompting
Prior to starting the work we tuned the prompt phrasing for best performance. When using $n$ exemplars, we used the first $n$ points from the development set. For AM we prompted GPT-4 with and without CoT. CoT conditions the LLM to work step-by-step by following a templatized process (e.g., “Let’s think about this step-by-step…”). It is known to provide good results in multiple reasoning tasks [8, 32, 15]. All our prompts followed the structure from GPT-4’s technical report [1], as we observed it produced more reliable outputs. For CoT we followed a tuned template (“Let’s read line-by-line and solve this step by step”) indicating which sentence was being read, as well as the rationale. For example, “We now read {SENTENCE}. It follows the previous argument, and hence it is labelled with an ‘I’.” Sample prompts can be found in Figures LABEL:fig:sampleamconcrete and 2(a).
RRv2 is measured with binary F1 (F01) for APE, and micro-F1 (Fμ) for AM [5]. Our prompts have specified a return format to signal the beginning of parsing.
3.4 Baselines
Our baselines are the MLMC [28] and MRC-APE [27] models. MLMC approaches APE as a table-filling problem: passages are related by their pairing on a table and it relies on an especially designed encoding scheme and loss. MRC-APE phrases it as a reading comprehension task: first, the model does AM, and then pairs the detected arguments. This approach is effective when using longer-context layers, of up to $4,096$ tokens. We additionally consider random guessers for AM (around 33% Fμ) and APE (14% F01).
@@ 3.4 Baselines
Extract from the passage all the arguments.
@@
In APE we ask for the pairing (e.g., return all arguments from the response that match those of the review”).
@@
Label the beginning of every argument with a
"B". Label the rest of the argument with an
"I". Label every line that is not part of an
argument with an "O".
@@
(a) Sample symbolic AM prompt. The model must return BIO tags. In other representations (e.g., indices) the model must to mark the B-label in parenthesis (e.g, “(15) 16 17”). In APE this output is of the form “argument lines: response lines”, and we convert into a binary matrix for scoring.
Figure 2: Sample prompts for our concrete (left) and symbolic (right) settings. Exemplars, if any, are included in the prompt. For zero-shot we only specify the output representation. Note how the actual task definition and prompt structure remains unchanged, and we only alter the I/O representations.
@@ 3.5 Settings
We have named our settings (i.e., representations) as concrete and symbolic, to distinguish the approach taken towards representing the task. Concrete returns full sentences, while symbolic encompasses a variety of I/O symbols. This is only for practical purposes: symbolic approaches cover multiple I/O representations, some of which may be easier than concrete; and, strictly speaking, the concrete setting is a type of symbolic representation [33]. See Table 1 for a full description of the settings tested.
The concrete setting we instruct the LLM to return lines in text based on the prompt: in AM, it must be part of an argument, in APE, an argument pair. To distinguish the “B” and “I” labels, we enforce a specific return format to work with our parsing code via a special token (|START|). For scoring concrete settings we expect an exact text match.
In symbolic settings the LLM must return symbols (labels) based on the prompt. This requires more reasoning steps than in concrete settings: the LLM is solving AM and labelling the span with an arbitrary label set defined in the prompt. In AM symbolic we evaluated two types of labels: BIO tags and line indices. For APE we evaluated line indices and the full binary matrix representation. We also used abstract meaning representation (AMR [34]) graphs, which are used in argument interpretation [35].
Task Input Output (label set) Abstraction
AM* Text Text and |START| Lowest
AM* Text with indices Indices Low
AM* Text Indices Medium
AM* Text BIO tags Medium-high
AM Text with AMR graph BIO tags High
AM AMR graph BIO tags Highest
Table 1: Input representations tested, in roughly increasing order of abstraction. Tasks marked with an asterisk (*) were tested with and without CoT. The first row is our concrete reasoning setting. Our ranking of abstraction is arbitrary: we consider the text with indices marked inline less abstract than a text without them, since the former provides a “hint” of what the label set is supposed to be like. Output representations with BIO tags as the output are more abstract, since it requires rule matching to determine the labeling.
@@ 4 Experiments and Results
We report our results comparing raw scores with respect to our settings (Section 4.1); number of exemplars (Section 4.2); and I/O representations (Section 4.3).
@@ 4.1 AM/APE: Symbolic and Concrete Reasoning
Results for the best-performing prompts and settings are in Table 2, and a description of every setting in Table 1. In AM the LLMs did well but did not beat the SOTA. The best-performing symbolic setting had line indices included in the input representation, and requested the indices of each argument as the output representation. To convert to BIO tags, we instructed the model to return the “B” labels as indices enclosed in parentheses. Not including the indices in the input did not lead to an equivalently good performance.
For APE, GPT-4 consistently bested the best-performing models ($+14\%$ F01). Both symbolic and concrete approaches did well with respect to the existing non-LLM-based baselines. We tested other approaches, such as first extracting the arguments and then matching them, but it did not yield sufficiently good results. Requesting a binary matrix output led to extremely poor performance ($9.88\%$ F01, below random). Due to token-length and budget limitations, we were unable to test CoT and AMR in APE.
CoT approaches had generally better performance than their non-CoT counterparts, even though they use fewer exemplars. The only exception to this was symbolic prompting (indices inline and indices as output) where the difference was $3\%$ points.
Model AM Fμ APE F01
GPT-3 (concrete) $39.86\pm 0.51$ $18.58\pm 0.70$
GPT-4 (concrete) $64.51\pm 0.53$ $53.84\pm 0.73$
GPT-3 (symbolic) $62.00\pm 0.32$ $20.15\pm 0.91$
GPT-4 (symbolic) $70.63\pm 0.21$ $49.85\pm 0.96$
@@
Table 2: Results for the AM and APE tasks in our settings. The best-performing symbolic prompt had indices inline and indices as the output. We also report MLMC and MRC-APE, the two best-performing, non-LLM-based approaches for RRv2. GPT-4 almost matched the existing baselines in AM and bested them in APE.